OpenStudy (anonymous):

A perfume bottle is a regular pyramid with square base. If the base has side 4 cm and a slant height 3cm, how many cubic centimeters of perfume will the bottle hold?

OpenStudy (dumbcow):

volume = 1/3*AreaBase*height AreaBase = 4^2 = 16 height = sqrt(3^2 - 2^2) = sqrt(5) volume = 1/3*16*sqrt(5)

OpenStudy (anonymous):

I get 11.9 cubic centimeters

OpenStudy (dumbcow):

yes correct